An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation fo $78.11 a)At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket? state your hypothesis and decision rule. b)Is this a close decision?
Answer by stanbon(75887)

An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation fo $78.11
-----------------------------------------------------------------------
a)At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket?
-----------------------------------------------------------------------
state your hypothesis and decision rule.
Ho: u <= 250
H1: u > 250
--------------------------------
test stat: t(275.66)= (275.66 - 250)/[78.11/sqrt(25)] = 1.6426
---
p-value: P(t> 1.6426 when df = 24) = tcdf(1.6425,100,24) = 0.0568
----------------------
Conclusion: Since the p-value is greater than 5%, fail to reject Ho.
At the 5% significance level the test results show the sample does
not violate the guideline.
------------------------
b)Is this a close decision?
Yes, since the P-value is close to 5%
=========================================
Cheers,
Stan H.